Your firm is 76% equity financed. You have a cost of debt of 3.8, a cost of equity of 18.7, and a tax rate of 17. What is your firm's weighted average cost of capital? {enter your number as a percentage to two decimal places.
Answer #1

D/A = 1-E/A = 1-0.76 = 0.24

After tax cost of debt = cost of debt*(1-tax rate)
After tax cost of debt = 3.8*(1-0.17)
= 3.154
WACC=after tax cost of debt*W(D)+cost of equity*W(E)
WACC=3.15*0.24+18.7*0.76
WACC =14.97%